Find the standard deviation of the expected value of the probability distribution, X={1, 2, 3, 4, 5} P(X)={.10, .35, .15, .23, .17}


1
Expert's answer
2022-03-05T07:31:09-0500

"\mu=1*0.10+2*0.35+3*0.15+4*0.23+5*0.17=3.02."


"\sigma=\sqrt{1^2*0.10+2^2*0.35+3^2*0.15+4^2*0.23+5^2*0.17-3.02^2}="

"=\sqrt{0.2996}\approx0.5474."